OpenStudy (anonymous):

what is the slope of the line passing through the following points? (-2,-1) and (8,-3)

OpenStudy (anonymous):

(-3 - (-1)) / (8-(-2)) = -4/10

OpenStudy (anonymous):

thanks

OpenStudy (aroub):

-3+1/ 8+2 = -2/10 = -1/5